# Privacy Risks and Preservation Methods in Explainable Artificial Intelligence: A Scoping Review

### v2

@@ 1.1 Paradigm shift in technology and the need for explanations
Traditional software development processes have metamorphosed into stable and reliable frameworks through decades of fine tuning by software experts. These software systems are built on human designed algorithms and produce a trace of the logic used to generate an output. Even in complex systems, it is possible for software experts to analyze the logic and generate an explanation for a specific result. During the software development lifecycle, engineers focus on creating the algorithm and validating using well designed test cases that closely replicate real world scenarios. In contrast, modern AI systems do not have an underlying human-written algorithm and learn from data fed to them. This data-driven nature creates dependence of the system on data quality (Merhi, 2022) and introduces problems such as lack of fairness when data is biased, or irrelevant results when data is incomplete or outdated (Trocin et al., 2021). During the AI development phase, engineers access training datasets which may contain personally identifiable or sensitive information about individuals. For neural network systems, the development process often involves a trial-and-error approach, where high accuracy is targeted by tweaking the hyperparameters such as the learning rate, epochs, number of layers or activation functions. The lack of an algorithm prevents engineers from tracing through the AI system and interpreting the results. Thus, the basic ability to be explainable and understand input-output behaviors, which is critical to all computer systems (Sundararajan et al., 2017), is often out of reach of AI systems. Explanations for outcomes of AI are crucial in high-risk applications (Mochaourab et al., 2023) in domains such as healthcare (Dhar et al., 2023; Dwivedi et al., 2023; Yang et al., 2022), finance (Dwivedi et al., 2023; Zhang et al., 2022), defense (Dwivedi et al., 2023), justice (Deeks, 2019), energy and power (Machlev et al., 2022) where the impact on human life and well-being is significant (Karimi et al., 2023; McDermid et al., 2021; Nassar et al., 2020) and the inability to do so deters their successful implementation (Nassar et al., 2020; Shrikumar et al., 2017a; Yang et al., 2022).
@@
Trustworthy AI strives to mitigate risks due to possible harms from the data-driven nature of AI systems. Trustworthiness is based on foundation principles of reliability, validity, robustness, privacy, explainability and fairness (Alzubaidi et al., 2023; Tabassi, 2023) to boost user confidence in the system outputs. Among these principles, explainability aims to bring the much-needed transparency in opaque models and can be considered as a non-functional requirement of a software system to mitigate opacity (Chazette et al., 2021). There are numerous benefits of including explanations in AI models. Besides aiding data scientists in getting a better understanding of the data (Hohman et al., 2019) and performing required data cleansing (Chen et al., 2022b), explanations can help developers in detecting errors in input and determining features that can be modified to change the outcome (Datta et al., 2016). When multiple models are available with similar accuracy, an explanation method can help to choose between models (Dhurandhar et al., 2018). Interpretable models can enable knowledge discovery by detecting knowledge or patterns that were missed by uninterpretable ones (Kim et al., 2016). Since humans remain an important component in the decision-making process as end-users and consumers of automated decisions (Terziyan & Vitko, 2022), explanations can give them an understanding of the model outcome, especially when they are adversely affected by the decisions (Ali et al., 2023). Explainability can also facilitate privacy awareness in end-users (Brunotte et al., 2021), enabling them to make right choices for their personal data and aid regulators and compliance officers to understand the compliance of models (McDermid et al., 2021) with applicable regulations. With generative AI (Gen-AI) and large language models (LLMs) entering mainstream, explanations constitute an important design principle (Weisz et al., 2023) in enabling a better mental model for users (Sun et al., 2022) and in communicating its capabilities and limitations to them (Weisz et al., 2023). It can also support users in effective prompt engineering to determine the words that impact the output of a model (Mishra et al., 2023) and in verification of generated content to mitigate the problem of hallucinations (Schneider, 2024b).
@@ 1.2 Challenges for privacy in explainability
In many high risk application domains of AI, training models on sensitive personal information is inevitable for usefulness of these systems (Veugen et al., 2022). For instance, a lung cancer detection model necessitates training on chest X-ray images, which constitutes personal information of patients. Similarly, a loan evaluation model of a bank, requires access to the financial profiles of customers, which is also personal information of individuals. Usage of personal data impacts the privacy of individuals when they are subject to intentional or unintentional identification and exposure through these systems. Some models are found to memorize data contained in the input (Song et al., 2017) which can be exploited by adversaries for extraction of personal information. Gen-AI models create new content from large multi-modal datasets (Sun et al., 2022) which could potentially contain sensitive personal information (Meskó & Topol, 2023). Due to such privacy risks involved, when personal data is used in training, testing, or inferencing of AI models, they become subject to data regulation and privacy acts (ICO, 2020).
Explainability is a foundational principle of Trustworthy AI, however, recent research has determined that introducing explanations in AI systems is found to conflict with the privacy requirements of the system. Explanation interfaces are found to give adversaries an additional attack surface (Duddu & Boutet, 2022; Liu et al., 2024) to mine the information contained in the model. Privacy attacks can target explanations to retrieve information about membership in the training set (Liu et al., 2024; Naretto et al., 2022; Shokri et al., 2021), build surrogates of the underlying model (Aïvodji et al., 2020; Wang et al., 2022; Yan et al., 2023b), infer sensitive attributes of individuals (Duddu & Boutet, 2022; Luo et al., 2022) and reconstruct the training set (Shokri et al., 2021). This leakage is demonstrated across different types of XAI methods including those that are currently used in commercial production systems. In addition to privacy attacks, the content of explanations may also inadvertently expose information that is proprietary (Milli et al., 2019) and hence valuable and confidential to organizations (Winikoff & Sardelic, 2021) or sensitive to individuals, thus causing breach of data and privacy regulations. Hence researchers have highlighted the urgent need of mitigating privacy leakage through explanation interfaces (Luo et al., 2022; Patel et al., 2022; Yan et al., 2023b). Due to these concerns of the privacy vulnerabilities of explanations, necessary privacy preservation measures are required in XAI systems (Aïvodji et al., 2020; Shokri et al., 2021; Zhao et al., 2021).
@@ 1.3 Main contributions
Previous research has identified that the privacy issues in explainability are insufficiently studied (Liu et al., 2024; Luo et al., 2022; Naretto et al., 2022) despite its criticality in achieving safety in AI transparency. To the best of our knowledge, there is currently no work that provides an in-depth understanding of the conflict between privacy and explainability in AI. Hence, we focus this article on these two fundamental desiderata of Trustworthy AI and explore the landscape of privacy risks and preservation methods proposed in literature in the context of XAI. The key questions that we have designed to define the scope of this article are:
@@ 2.1 Definition of XAI
In 2017, DARPA kickstarted its 4-year XAI program to accelerate research in the development of explanation methods and interfaces to enhance understanding and trust of end-users (Gunning & Aha, 2019). The program defined XAI as “AI systems that can explain their rationale to a human user, characterize their strengths and weaknesses, and convey an understanding of how they will behave in the future” (Gunning & Aha, 2019). The study established users’ preference for systems with explanations over systems that provided only decisions. Ribeiro et al. (2016) refer to explanations of predictions as qualitative artifacts that provide the relationship between an input instance and the output prediction.
@@ 2.2 Evolution of XAI and emergence of privacy concerns
The field of explainability can be traced to the early 1990s, driven by the lack of transparency in black-box models. Early contributions (Benitez et al., 1997; Craven & Shavlik, 1995; LiMin Fu, 1994; Milaré et al., 2002; Torres & Rocco, 2005) proposed different techniques for extracting interpretable representations from these systems. The rise of deep learning and the improvement in the predictive performance of black-box systems, propelled complex uninterpretable systems into mainstream usage. However, their use in critical domains remains problematic due to their lack of transparency. Regulatory frameworks such as the General Data Protection Regulation (GDPR, 2016), specifically the provisions on individuals’ rights related to automated decision-making including profiling, intensified the demand for transparent, explainable models thus resulting in a rapid growth in the field of explainability.
However, the introduction of transparency through XAI methods has also exposed new vectors for privacy leakage through explanation interfaces. Early studies (Milli et al., 2019; Shokri et al., 2020; 2021) described privacy attacks on the training data and the underlying model. In response, researchers have begun to explore defense mechanisms and pioneering works in this field (Harder et al., 2020; Patel et al., 2022) have proposed various strategies for generating privacy preserved explanations. Despite these efforts, privacy risks in XAI remain an open research problem, with novel attacks being identified and defense strategies being actively investigated. Figure 1 outlines the key milestones in the evolution of XAI and highlights the emergence of privacy issues and proposed defenses.
@@ 2.3 Categorization of XAI
In recent years, several different XAI methods have been proposed. Broadly, explainability can be achieved using inherently interpretable models or applying post-hoc methods on trained models (Harder et al., 2020). Methods specific to certain model types and capabilities, are referred to as model-specific while those independent of the model are referred to as model-agnostic (Dwivedi et al., 2023). In this subsection, we discuss the main categories into which XAI methods are grouped in existing literature (Table LABEL:tab:table1), based on the underlying mechanism used to derive explanations. We also include categories suitable for Gen-AI (Sections  2.3.5 to  2.3.8) that explore new explanation paradigms for large models (Sarkar et al., 2024). Since there is a broad spectrum of available explainability methods, we limit ourselves to a selection of methods to give readers sufficient understanding of the terminologies used in subsequent sections. For a comprehensive review of XAI categories and methods, we refer the reader to other related reviews listed in Section 2.4.
@@ 2.3.1 Interpretable methods
These AI models are understandable by design (Arrieta et al., 2020). They have embedded rules or transparent architecture that facilitates the understanding of the input-output logic of the system. They are also referred to as white-box or transparent models. Decision trees, Bayesian networks, linear/logistic regression, k-nearest neighbours, rule based systems and general additive models (Arrieta et al., 2020; Molnar, 2023; Rawal et al., 2022) are some examples of interpretable models. Though these models are promising in aiding the understandability of a system, they have limitations. The primary deterrent to their successful adoption as explainable-by-design methods, is their lower accuracy (Blanco-Justicia et al., 2020; El Zein et al., 2024; Gunning & Aha, 2019) compared to better performing black-box models such as deep learning systems. They also lack natural language explanations, making them unsuitable for use by non-technical users (Biran & McKeown, 2017). Nonetheless, due to their intrinsically transparent architecture, interpretable models are often used as surrogates for black-box models (McDermid et al., 2021). The use of multiple surrogate models is found to facilitate the availability of different types of explanations (Dwivedi et al., 2023) improving the overall interpretability of the system.
@@ 2.3.2 Example-based methods
These methods use examples, i.e., data instances as samples to explain the model (McDermid et al., 2021). The instances may be from the training set or generated by the method (Jiménez-Luna et al., 2020; Li et al., 2020b). These methods are also referred to as record-based (Shokri et al., 2020), instance-based (Jiménez-Luna et al., 2020) or case-based (Montenegro et al., 2021) methods in literature. They can complement feature-based methods to aid understandability of the end users (Jia et al., 2021) and also improve the interpretability of complex distributions (Kim et al., 2016). They are intuitive and natural in their ability to provide explanations to humans (Jiménez-Luna et al., 2020). Some methods in this category are anchors (Ribeiro et al., 2018), contrastive explanations (Dhurandhar et al., 2018), counterfactuals (Wachter et al., 2017), influence functions (Koh & Liang, 2017) and prototypes and criticisms (Kim et al., 2016).
@@ 2.3.3 Knowledge-based methods
These methods utilize knowledge representation techniques in machine learning (ML) models to enhance interpretability (Tiddi & Schlobach, 2022). The integration of background knowledge (Hitzler et al., 2022) facilitates the incorporation of contextual information (Lecue, 2020; Páez, 2019), thus increasing the trustworthiness of explanations. The emerging field of neuro-symbolic (Hitzler et al., 2022) or in-between methods (Ilkou & Koutraki, 2020) explores the integration of symbolic AI approaches rooted in knowledge representation and reasoning with subsymbolic or connectionist based approaches (Hitzler et al., 2020).
Another knowledge-based approach is the use of semantic web technologies for semantic interpretation and automated reasoning from structured knowledge bases (Seeliger et al., 2019). Knowledge graphs and ontologies are the common tools that can be deployed to support explainability. Knowledge graphs have applicability in pre-model and post-model explainabilty contexts for feature extraction, relation identification, inferencing and reasoning (Rajabi & Etminani, 2022). The field of semantic web technologies in explainability is attractive because of its potential in creating knowledge-rich explanations without compromising the model performance (Seeliger et al., 2019).

### v3

@@ 1.1 Paradigm shift in technology and the need for explanations
Traditional software development processes have metamorphosed into stable and reliable frameworks through decades of fine tuning by software experts. These software systems are built on human designed algorithms and produce a trace of the logic used to generate an output. Even in complex systems, it is possible for software experts to analyze the logic and generate an explanation for a specific result. During the software development lifecycle, engineers focus on creating the algorithm and validating using well designed test cases that closely replicate real world scenarios. In contrast, modern AI systems do not have an underlying human-written algorithm and learn from data fed to them. This data-driven nature creates dependence of the system on data quality (merhiAssessmentBarriersImpacting2022) and introduces problems such as lack of fairness when data is biased, or irrelevant results when data is incomplete or outdated (trocinResponsibleAIDigital2021). During the AI development phase, engineers access training datasets which may contain personally identifiable or sensitive information about individuals. For neural network systems, the development process often involves a trial-and-error approach, where high accuracy is targeted by tweaking the hyperparameters such as the learning rate, epochs, number of layers or activation functions. The lack of an algorithm prevents engineers from tracing through the AI system and interpreting the results. Thus, the basic ability to be explainable and understand input-output behaviors, which is critical to all computer systems (sundararajanAxiomaticAttributionDeep2017), is often out of reach of AI systems. Explanations for outcomes of AI are crucial in high-risk applications (mochaourabDemonstratorCounterfactualExplanations2023) in domains such as healthcare (dharChallengesDeepLearning2023; dwivediExplainableAIXAI2023; yangUnboxBlackboxMedical2022), finance (dwivediExplainableAIXAI2023; zhangExplainableArtificialIntelligence2022), defense (dwivediExplainableAIXAI2023), justice (deeksJudicialDemandExplainable2019), energy and power (machlevExplainableArtificialIntelligence2022) where the impact on human life and well-being is significant (karimiSurveyAlgorithmicRecourse2023; mcdermidArtificialIntelligenceExplainability2021; nassarBlockchainExplainableTrustworthy2020) and the inability to do so deters their successful implementation (nassarBlockchainExplainableTrustworthy2020; shrikumarLearningImportantFeatures2017; yangUnboxBlackboxMedical2022).
@@
Trustworthy AI strives to mitigate risks due to possible harms from the data-driven nature of AI systems. Trustworthiness is based on foundation principles of reliability, validity, robustness, privacy, explainability and fairness (alzubaidiRiskFreeTrustworthyArtificial2023; tabassiAIRiskManagement2023) to boost user confidence in the system outputs. Among these principles, explainability aims to bring the much-needed transparency in opaque models and can be considered as a non-functional requirement of a software system to mitigate opacity (chazetteExploringExplainabilityDefinition2021). There are numerous benefits of including explanations in AI models. Besides aiding data scientists in getting a better understanding of the data (hohmanGamutDesignProbe2019) and performing required data cleansing (chenEVFLExplainableVertical2022), explanations can help developers in detecting errors in input and determining features that can be modified to change the outcome (dattaAlgorithmicTransparencyQuantitative2016). When multiple models are available with similar accuracy, an explanation method can help to choose between models (dhurandharExplanationsBasedMissing2018). Interpretable models can enable knowledge discovery by detecting knowledge or patterns that were missed by uninterpretable ones (kimExamplesAreNot2016). Since humans remain an important component in the decision-making process as end-users and consumers of automated decisions (terziyanExplainableAIIndustry2022), explanations can give them an understanding of the model outcome, especially when they are adversely affected by the decisions (aliExplainableArtificialIntelligence2023). Explainability can also facilitate privacy awareness in end-users (brunotteCanExplanationsSupport2021), enabling them to make right choices for their personal data and aid regulators and compliance officers to understand the compliance of models (mcdermidArtificialIntelligenceExplainability2021) with applicable regulations. With generative AI (Gen-AI) and large language models (LLMs) entering mainstream, explanations constitute an important design principle (weiszGeneralDesignPrinciples2023) in enabling a better mental model for users (sunInvestigatingExplainabilityGenerative2022) and in communicating its capabilities and limitations to them (weiszGeneralDesignPrinciples2023). It can also support users in effective prompt engineering to determine the words that impact the output of a model (mishraPromptAidPromptExploration2023) and in verification of generated content to mitigate the problem of hallucinations (schneiderExplainableGenerativeAI2024).
@@ 1.2 Challenges for privacy in explainability
In many high risk application domains of AI, training models on sensitive personal information is inevitable for usefulness of these systems (veugenPrivacyPreservingContrastiveExplanations2022). For instance, a lung cancer detection model necessitates training on chest X-ray images, which constitutes personal information of patients. Similarly, a loan evaluation model of a bank, requires access to the financial profiles of customers, which is also personal information of individuals. Usage of personal data impacts the privacy of individuals when they are subject to intentional or unintentional identification and exposure through these systems. Some models are found to memorize data contained in the input (songMachineLearningModels2017) which can be exploited by adversaries for extraction of personal information. Gen-AI models create new content from large multi-modal datasets (sunInvestigatingExplainabilityGenerative2022) which could potentially contain sensitive personal information (meskoImperativeRegulatoryOversight2023). Due to such privacy risks involved, when personal data is used in training, testing, or inferencing of AI models, they become subject to data regulation and privacy acts (icoExplainingDecisionsMade2020).
Explainability is a foundational principle of Trustworthy AI, however, recent research has determined that introducing explanations in AI systems is found to conflict with the privacy requirements of the system. Explanation interfaces are found to give adversaries an additional attack surface (dudduInferringSensitiveAttributes2022; liuPleaseTellMe2024) to mine the information contained in the model. Privacy attacks can target explanations to retrieve information about membership in the training set (liuPleaseTellMe2024; narettoEvaluatingPrivacyExposure2022; shokriPrivacyRisksModel2021), build surrogates of the underlying model (aivodjiModelExtractionCounterfactual2020; wangDualCFEfficientModel2022; yanExplanationLeaksExplanationguided2023), infer sensitive attributes of individuals (dudduInferringSensitiveAttributes2022; luoFeatureInferenceAttack2022) and reconstruct the training set (shokriPrivacyRisksModel2021). This leakage is demonstrated across different types of XAI methods including those that are currently used in commercial production systems. In addition to privacy attacks, the content of explanations may also inadvertently expose information that is proprietary (milliModelReconstructionModel2019) and hence valuable and confidential to organizations (winikoffArtificialIntelligenceRight2021) or sensitive to individuals, thus causing breach of data and privacy regulations. Hence researchers have highlighted the urgent need of mitigating privacy leakage through explanation interfaces (luoFeatureInferenceAttack2022; patelModelExplanationsDifferential2022; yanExplanationLeaksExplanationguided2023). Due to these concerns of the privacy vulnerabilities of explanations, necessary privacy preservation measures are required in XAI systems (aivodjiModelExtractionCounterfactual2020; shokriPrivacyRisksModel2021; zhaoExploitingExplanationsModel2021).
@@ 1.3 Main contributions
Previous research has identified that the privacy issues in explainability are insufficiently studied (liuPleaseTellMe2024; luoFeatureInferenceAttack2022; narettoEvaluatingPrivacyExposure2022) despite its criticality in achieving safety in AI transparency. To the best of our knowledge, there is currently no work that provides an in-depth understanding of the conflict between privacy and explainability in AI. Hence, we focus this article on these two fundamental desiderata of Trustworthy AI and explore the landscape of privacy risks and preservation methods proposed in literature in the context of XAI. The key questions that we have designed to define the scope of this article are:
@@ 2.1 Definition of XAI
In 2017, DARPA kickstarted its 4-year XAI program to accelerate research in the development of explanation methods and interfaces to enhance understanding and trust of end-users (gunningDARPAExplainableArtificial2019). The program defined XAI as “AI systems that can explain their rationale to a human user, characterize their strengths and weaknesses, and convey an understanding of how they will behave in the future” (gunningDARPAExplainableArtificial2019). The study established users’ preference for systems with explanations over systems that provided only decisions. ribeiroWhyShouldTrust2016 refer to explanations of predictions as qualitative artifacts that provide the relationship between an input instance and the output prediction.
@@ 2.2 Evolution of XAI and emergence of privacy concerns
The field of explainability can be traced to the early 1990s, driven by the lack of transparency in black-box models. Early contributions (benitezAreArtificialNeural1997; cravenExtractingTreeStructuredRepresentations1995; liminfuRuleGenerationNeural1994; milareApproachExplainNeural2002; torresExtractingTreesTrained2005) proposed different techniques for extracting interpretable representations from these systems. The rise of deep learning and the improvement in the predictive performance of black-box systems, propelled complex uninterpretable systems into mainstream usage. However, their use in critical domains remains problematic due to their lack of transparency. Regulatory frameworks such as the General Data Protection Regulation (gdprArt22GDPR2016), specifically the provisions on individuals’ rights related to automated decision-making including profiling, intensified the demand for transparent, explainable models thus resulting in a rapid growth in the field of explainability.
However, the introduction of transparency through XAI methods has also exposed new vectors for privacy leakage through explanation interfaces. Early studies (milliModelReconstructionModel2019; shokriExploitingTransparencyMeasures2020; shokriPrivacyRisksModel2021) described privacy attacks on the training data and the underlying model. In response, researchers have begun to explore defense mechanisms and pioneering works in this field (harderInterpretableDifferentiallyPrivate2020; patelModelExplanationsDifferential2022) have proposed various strategies for generating privacy preserved explanations. Despite these efforts, privacy risks in XAI remain an open research problem, with novel attacks being identified and defense strategies being actively investigated. Figure 1 outlines the key milestones in the evolution of XAI and highlights the emergence of privacy issues and proposed defenses.
@@ 2.3 Categorization of XAI
In recent years, several different XAI methods have been proposed. Broadly, explainability can be achieved using inherently interpretable models or applying post-hoc methods on trained models (harderInterpretableDifferentiallyPrivate2020). Methods specific to certain model types and capabilities, are referred to as model-specific while those independent of the model are referred to as model-agnostic (dwivediExplainableAIXAI2023). In this subsection, we discuss the main categories into which XAI methods are grouped in existing literature (Table LABEL:tab:table1), based on the underlying mechanism used to derive explanations. We also include categories suitable for Gen-AI (Sections  2.3.5 to  2.3.8) that explore new explanation paradigms for large models (sarkar2024explaining). Since there is a broad spectrum of available explainability methods, we limit ourselves to a selection of methods to give readers sufficient understanding of the terminologies used in subsequent sections. For a comprehensive review of XAI categories and methods, we refer the reader to other related reviews listed in Section 2.4.
@@ 2.3.1 Interpretable methods
These AI models are understandable by design (arrietaExplainableArtificialIntelligence2020). They have embedded rules or transparent architecture that facilitates the understanding of the input-output logic of the system. They are also referred to as white-box or transparent models. Decision trees, Bayesian networks, linear/logistic regression, k-nearest neighbours, rule based systems and general additive models (arrietaExplainableArtificialIntelligence2020; molnarInterpretableMachineLearning2023; rawalRecentAdvancesTrustworthy2022) are some examples of interpretable models. Though these models are promising in aiding the understandability of a system, they have limitations. The primary deterrent to their successful adoption as explainable-by-design methods, is their lower accuracy (blanco-justiciaMachineLearningExplainability2020; elzeinPrivaTreeCollaborativePrivacyPreserving2024; gunningDARPAExplainableArtificial2019) compared to better performing black-box models such as deep learning systems. They also lack natural language explanations, making them unsuitable for use by non-technical users (biranHumanCentricJustificationMachine2017). Nonetheless, due to their intrinsically transparent architecture, interpretable models are often used as surrogates for black-box models (mcdermidArtificialIntelligenceExplainability2021). The use of multiple surrogate models is found to facilitate the availability of different types of explanations (dwivediExplainableAIXAI2023) improving the overall interpretability of the system.
@@ 2.3.2 Example-based methods
These methods use examples, i.e., data instances as samples to explain the model (mcdermidArtificialIntelligenceExplainability2021). The instances may be from the training set or generated by the method (jimenez-lunaDrugDiscoveryExplainable2020; liSurveyDatadrivenKnowledgeaware2020). These methods are also referred to as record-based (shokriExploitingTransparencyMeasures2020), instance-based (jimenez-lunaDrugDiscoveryExplainable2020) or case-based (montenegroPrivacyPreservingGenerativeAdversarial2021) methods in literature. They can complement feature-based methods to aid understandability of the end users (jiaPredictionWeaningMechanical2021) and also improve the interpretability of complex distributions (kimExamplesAreNot2016). They are intuitive and natural in their ability to provide explanations to humans (jimenez-lunaDrugDiscoveryExplainable2020). Some methods in this category are anchors (ribeiroAnchorsHighPrecisionModelAgnostic2018), contrastive explanations (dhurandharExplanationsBasedMissing2018), counterfactuals (wachterCounterfactualExplanationsOpening2017), influence functions (kohUnderstandingBlackboxPredictions2017) and prototypes and criticisms (kimExamplesAreNot2016).
@@ 2.3.3 Knowledge-based methods
These methods utilize knowledge representation techniques in machine learning (ML) models to enhance interpretability (tiddiKnowledgeGraphsTools2022). The integration of background knowledge (hitzlerNeurosymbolicApproachesArtificial2022) facilitates the incorporation of contextual information (lecueRoleKnowledgeGraphs2020; paezPragmaticTurnExplainable2019), thus increasing the trustworthiness of explanations. The emerging field of neuro-symbolic (hitzlerNeurosymbolicApproachesArtificial2022) or in-between methods (ilkouSymbolicVsSubsymbolic2020) explores the integration of symbolic AI approaches rooted in knowledge representation and reasoning with subsymbolic or connectionist based approaches (hitzlerNeuralsymbolicIntegrationSemantic2020).
Another knowledge-based approach is the use of semantic web technologies for semantic interpretation and automated reasoning from structured knowledge bases (seeligerSemanticWebTechnologies2019). Knowledge graphs and ontologies are the common tools that can be deployed to support explainability. Knowledge graphs have applicability in pre-model and post-model explainabilty contexts for feature extraction, relation identification, inferencing and reasoning (rajabiKnowledgegraphbasedExplainableAI2022). The field of semantic web technologies in explainability is attractive because of its potential in creating knowledge-rich explanations without compromising the model performance (seeligerSemanticWebTechnologies2019).